# On the polarimetric response of the Nançay Radio Telescope and its impact on precision pulsar timing

### v1

@@ Abstract
Context. Precision pulsar timing and studies of pulsar radio emission properties require accurate polarimetric calibration of the radio observations, since incorrect calibration can distort profiles and introduce noise in Time of Arrival (TOA) data. In Guillemot et al. (2023) we presented a new method for calibrating pulsar observations conducted with the Nançay decimetric Radio Telescope (NRT), which significantly improved NRT polarimetric measurements and pulsar timing quality for data taken after this method was developed, in November 2019.
Aims. Results from the above-mentioned study hinted at a dependence of the polarimetric response of the NRT on the observed direction. We therefore investigated this potential dependence, since unaccounted variations of the instrumental response could degrade polarimetric measurements. Additionally, we aimed to develop a method for properly calibrating NRT pulsar observations conducted before November 2019.
Methods. We conducted three series of observations of bright pulsars over wide declination ranges, in a special observation mode in which the feed horn rotates by 180∘ degrees across the observation, enabling us to determine the full polarimetric response of the NRT while modeling potential variations of calibration parameters with hour angle and declination. In addition, we developed a method that uses the Measurement Equation Template Matching (METM) technique to improve the calibration of pre-November 2019 data.
Results. From the analysis of the series of observations of bright pulsars with horn rotation, we found that the polarimetric response of the NRT does not appear to vary with hour angle or declination. On the other hand, the new METM-based calibration method appears to significantly improve the calibration of pre-November 2019 data. By analyzing NRT data on a selection of millisecond pulsars we found that the new polarimetric profiles are more homogeneous, they generally have larger signal-to-noise ratios, and found that the TOA data for these MSPs are more accurate and contain lower levels of noise, especially when combining the new calibration method with the Matrix Template Matching (MTM) method for extracting TOAs from pulsar observations.
@@
Pulsars are rotation-powered neutron stars that emit beams of electromagnetic radiation. These beams are swept across the sky with the rotation of the stars, so that distant observers whose lines of sight are crossed by the beams might detect periodic pulses. The pulsar timing technique consists in modeling the rotation of pulsars by measuring the Times of Arrival (TOAs) of the pulses detected at a telescope and comparing them to the predictions from a timing model. Differences between measured and predicted TOAs, the so-called timing residuals, contain information about the pulsars’ spin properties, orbital properties (for pulsars in multiple systems), astrometric properties, or information about the interstellar medium or their immediate environment (see e.g. Chapter 2 of Lorimer & Kramer, 2004, for a complete description). Applications of pulsar timing are numerous and diverse: for example, tests of General Relativity (GR) in the strong-field regime (Kramer et al., 2021), searches for exoplanets orbiting pulsars (Niţu et al., 2024; Voisin et al., 2024), measurements of pulsar masses and constraints on the equations-of-state of neutron star masses using these measurements (e.g., Özel & Freire, 2016), or searches for low-frequency gravitational waves using Pulsar Timing Arrays (PTAs; see FERMI-LAT Collaboration et al., 2022; Agazie et al., 2023; Reardon et al., 2023; EPTA Collaboration et al., 2023b, for recent results from the different PTA collaborations).
@@
In standard radio pulsar timing analyses, TOAs for a given pulsar are determined by comparing radio observations to a standard profile, assumed to be a noise-free representation of the pulsar’s intrinsic pulse profile shape. The standard profiles can be used to determine TOAs for arbitrarily large fractions of the recorded frequency bandwidths, assuming that the pulse profile does not evolve with frequency within these frequency intervals. Alternatively, the standard profiles can encode frequency evolution of their component(s), leading to more accurate TOAs in the case of large bandwidth observations (the latter wide-band template matching technique is for instance implemented in the PulsePortraiture software library, see Pennucci et al., 2014). Independently of the method used for constructing the template profiles, standard pulsar timing analyses crucially rely on the assumption that the template profiles are faithful representations of the pulsar pulse profiles. Reciprocally, individual observations of a given pulsar must resemble the standard profile and not vary; if not, the cross-correlation between the individual observations and the standard profile will lead to inaccurate TOAs and introduce systematic timing errors in the TOAs.
@@
If not properly accounted for, pulse profile variations can cause individual observations of a pulsar to differ from the chosen standard profile. In some cases, these variations originate from the pulsar itself, e.g. from pulse jitter, moding, intermittency, or relativistic spin precession (see Padmanabh et al., 2021, and references therein). Profile variations can also result from instrumental effects. In particular, inaccurate polarization calibration can distort recorded pulse profiles and in turn introduce noise in TOA datasets (see for instance van Straten, 2013; Foster et al., 2015; Guillemot et al., 2023; Rogers et al., 2024). Such timing noise resulting from instrumental artefacts can degrade pulsar timing analyses, bias parameter measurements or mimic signals expected from stochastic gravitational wave background emission (Lentati et al., 2016).
@@
Accurate polarization calibration is therefore crucial for any pulsar timing studies, as well as for studying pulsar polarization properties. Recently, Dey et al. (2024) investigated different methods for calibrating pulsar observations done with the Green Bank Telescope, finding that the simplest method, the ideal feed assumption (hereafter IFA; Manchester et al., 2013) which assumes perfectly linear and orthogonal polarization feeds, led to better pulsar timing results than the more sophisticated Measurement Equation Modeling (MEM, van Straten, 2004) and Measurement Equation Template Matching (METM, van Straten, 2013) methods in which polarimetric responses are respectively determined from the analysis of strongly linearly polarized pulsars observed over wide ranges of parallactic angles, or by comparing observations of a given reference pulsar with a well-calibrated template profile for the same pulsar. Dey et al. (2024) attributed the poorer performances of the other tested methods to potential instabilities of the local polarization reference source, a noise diode coupled to the receiver, and time instabilities in the pulse profile of the pulsar used as reference for the METM method. The IFA, MEM and METM calibration methods are described in more detail later in this paper. Rogers et al. (2024), on the other hand, compared polarization calibration methods using pulsar data from the Murriyang 64m CSIRO Parkes radio telescope, and found that the METM method significantly improved pulsar timing quality over the IFA calibration scheme. Furthermore, Rogers et al. (2024) found that combining the METM data calibration method with the Matrix Template Matching (MTM) technique (van Straten, 2006), which uses all four Stokes components of the pulse profile and template profile when forming TOAs to compensate for residual inconsistencies in the polarimetric profiles, leads to significantly improved pulsar timing residuals. In their conclusions, Rogers et al. (2024) recommended that MTM be used for determining TOAs in PTA experiments. Finally, in previous work (see Guillemot et al., 2023) we significantly improved the polarization calibration of pulsar observations made with the Nançay decimetric Radio Telescope (NRT) in France, by determining full polarimetric responses of the instrument from the analysis of observations of a highly linearly polarized pulsar in a special observing mode where the feed horn rotates across the observation, and employing the MEM calibration model (details about this analysis are given in Sect. 2). The conclusion from this work is that the MEM-based calibration scheme outperforms IFA calibration for the NRT, with much improved polarimetric measurements, contrasting with the conclusions of Dey et al. (2024). Besides, similarly to Rogers et al. (2024) we also found that determining TOAs using the MTM technique significantly improved NRT pulsar timing quality, also suggesting that this TOA extraction technique should be used in high-precision pulsar timing stusdies.
@@
In this article, we present new work done following Guillemot et al. (2023) to further improve the polarization calibration of pulsar data taken with the NRT. The results from the observations of a bright, highly linearly polarized pulsar with feed horn rotation presented in Guillemot et al. (2023) hinted at an hour angle-dependence of the response of the NRT. We therefore conducted an analysis aimed at determining whether the instrumental response of the NRT depends on the observed direction, and more specifically, the hour angle and declination of the observed pulsar. In Sect. 3, we examine the possibility of properly-calibrating NRT data taken before the first observations of bright, highly polarized pulsars with horn rotation, in November 2019. This work, that used the METM calibration technique, led to the construction of a time- and frequency-dependent calibration model, that we applied to observations of millisecond pulsars (MSPs) carried out with the NRT. In Sect. 4 we present the new polarimetric profiles for these MSPs as well as an assessment of pulsar timing quality with the calibration scheme. We finally present a summary of our work in Sect. 5.
2 Does the polarimetric response of the NRT depend on the observed direction?
@@
The NRT is a meridian transit-type telescope of the Kraus/Ohio State design (see, for instance, Kraus, 1960, 1966) equivalent to a parabolic dish with a diameter of approximately 94m, and capable of observing objects with declinations above . Since August 2011, NRT pulsar observations are conducted with the NUPPI backend which uses Graphics Processing Units (GPUs) to coherently dedisperse and fold the dual linear polarization signals from the receiver in real time, over a total bandwidth of 512 MHz (see e.g. Desvignes et al., 2011; Cognard et al., 2013, for additional details on NUPPI). Pulsar observations with the NRT are primarily conducted at the central frequency of 1484 MHz. Further details on the NRT and on pulsar observations with the NUPPI backend can be found in Guillemot et al. (2023). In this new article, we will consider pulsar observations made with the NUPPI data at 1.4 GHz, leaving analyses of observations at other frequencies or observations conducted with the previous backend, the Berkeley-Orléans-Nançay (BON) pulsar backend, for future work. For all NUPPI data analyses presented in this new article we made use of the PSRCHIVE software library (Hotan et al., 2004), and cleaned the data of radio-frequency interference (RFI) using the Surgical method of the CoastGuard pulsar data analysis library (Lazarus et al., 2016).
@@
In the first article of this series (Guillemot et al., 2023), we presented an improved procedure for calibrating pulsar observations conducted with the NRT, developed in late 2019. Owing to its design, the NRT can track objects for, typically, one to two hours around transit (the maximum allowed tracking time depends on the observed declination). As a consequence, the NRT cannot sample wide ranges of parallactic angles when observing a given object. This limitation, in principle, prevents measurements of the full polarimetric response of the telescope with long observations of bright, highly polarized pulsars over wide ranges of parallactic angles (see e.g. Stinebring et al., 1984; Johnston, 2002; van Straten, 2004, for examples of such experiments). Nevertheless, in November 2019 we circumvented this limitation by beginning to carry out regular observations of the bright, highly linearly polarized pulsar J07422822 (also known as B074028, see Bonsignori-Facondi et al., 1973), in a non-standard observation mode in which the feed horn rotates by over the course of the 1 h observation. During these observations, the feed horn rotation effectively mimics wide parallactic angle variations, enabling much wider ranges of pulsar polarization angles to be sampled over the hour of observation. In turn, this enabled us to determine the full polarimetric response of the telescope. The calibration solutions obtained from these rotating horn observations enabled us to obtain vastly improved polarimetric results for the selection of millisecond pulsars (MSPs) that was used in our tests, with increased signal-to-noise ratios (S/N) in most cases and polarimetric profiles that were consistent with published results, unlike those obtained with the previously-used, simplistic calibration method.
@@
We began by investigating our analysis codes, and found an issue with the calculation of the time-varying parallactic for the NRT, in the analysis that led to the results presented in Fig. 9 of Guillemot et al. (2023). In the case of the NRT, the parallactic angle is given by:
@@
where denotes the declination of the observed source, HA is the hour angle of the source at a given time, and is the parallactic angle. The above expression is the same as the one given in Guillemot et al. (2023), and was the expression implemented in PSRCHIVE. However, we found that, due to the conventions used in PSRCHIVE for the signs and orientations of angles, the parallactic angle calculated using Eq. 1 needed to be used with the opposite sign in the software. We found that correcting for the sign of for NRT observations in PSRCHIVE significantly improved the results of the analysis of rotating horn observations. In Fig. 1 we show measured and modeled Stokes parameters for an observation of J07422822 with horn rotation conducted on MJD 59368 (3 June 2021), before and after correction. As can be seen from the right-hand panel of Fig. 1, measured and modeled values of the Stokes parameter in the first half of the observation are much more consistent. The fit results shown in the left-hand panel of Fig. 1 (i.e., before correction) give a reduced chi-square statistic of 2.25, while those shown in the right-hand panel give 1.73, indicating a better agreement between the data and the model. The analysis of all other rotating horn observations of J07422822 also found a better agreement between the data and the model, after the sign correction. In addition to this correction we made other improvements to the analysis pipeline. First, variations of the ionospheric Rotation Measure (RM) are now taken into account, using RMextract (Mevius, 2018) to determine ionospheric RM values for each sub-integration. Second, we now make use the “-Q” option of pcm to indicate that the reference noise diode is coupled after the frontend, as is indeed the case for our observations. We note, however, that the latter two modifications of the analysis pipeline compared to Guillemot et al. (2023) resulted in minor improvements to the analysis results.
@@
Beyond the correction made to the calculation of NRT parallactic angles in PSRCHIVE that significantly improved the agreement between measured and modeled Stokes parameters for rotating horn observations, the initial goal of the study was to test for potential dependencies of the response of the NRT on observation parameters. As was noted in Guillemot et al. (2023), the absolute gain of the NRT is known to vary with the declination of the observed object (see e.g. Theureau et al., 2005). To test for a dependence of calibration parameters on declination, and also test for a dependence on hour angle (e.g. resulting from the unusual design of the telescope), we therefore conducted observations of pulsars other than J07422822 in the special observation mode. We conducted a first series of rotating horn observations of pulsars at different declinations between MJDs 59368 and 59373 (3 and 8 June 2021). During this first series of observations, we sampled pulsar declinations between (J07422822) and (J0454+5543, a.k.a. B0450+55, Damashek et al., 1978). The other pulsars observed during this session were PSRs J0826+2637 (B0823+26, Craft et al., 1968), J0953+0755 (B0950+08, Pilkington et al., 1968), J1239+2453 (B1237+25, Lang, 1969), J17051906 (B170219, Manchester et al., 1978), and J2113+4644 (B2111+46, Davies & Large, 1970). The observations were conducted close in time (in this case, within five days), so that we could assume the instrumental response to be similar at all epochs. Table 1 lists the start epochs, declinations, durations, and start and end hour angles of these observations. Figs. 2 and 3 show the results of the MEM analysis of the individual observations and preceding observations of the reference noise diode. As in Fig. 1, left-hand (resp., right-hand) panels show the results obtained before (resp., after) correction of the sign of the NRT parallactic angle. Interestingly, inconsistencies between pulsars for some parameters (particularly, the differential gain , the feed orientation parameter and the Stokes parameter of the reference signal, i.e., parameters that constrain the polarization position angle, which strongly depends on the parallactic angle correction), appear to be significantly reduced by the correction. After correction, calibration and reference signal parameters seem consistent between pulsar observations, with the exception of the absolute gain , known to vary with declination.
@@
Strong RFI conditions can corrupt some of the observations, e.g. leading to entire frequency channels being blanked by RFI excision in these observations. If other observations are less affected then the analysis will determine the polarimetric response in more frequency channels than that of the individual, RFI-corrupted observations.
@@
Prior to this work, the pcm tool of PSRCHIVE did not allow multiple pulsar observations to be analyzed jointly. Therefore, in order to be able to conduct analyses of this type, we added new functionalities to pcm111Currently available under the “npsr-pcm” branch of PSRCHIVE, see git://git.code.sf.net/p/psrchive/code..
@@
In the above expressions, represents the 2D polynomial, is the degree of the polynomial in and is the degree of the polynomial in , and the terms represent coefficients to be fitted for. In our analysis, and respectively corresponded to the pulsar declination , and to the hour angle HA at a given time. We tested polynomial orders ranging from 0 (i.e., constant values) to 3 in HA, and ranging from 0 to 6 in declination. We used the same polynomial orders for both the differential gain and phase functions, to limit the number of polynomial order combinations to be tested. In the rest of the article we will denote the different analyses we performed using the following nomenclature: “A/B”, meaning that the 2D polynomials representing the differential gain and phase were of order A in declination, and of order B in hour angle.

### v2

# The polarimetric response of the Nançay Radio Telescope and its impact on precision pulsar timing

@@ Abstract
Context. Precision pulsar timing and studies of pulsar radio emission properties require accurate polarimetric calibration of the radio observations since incorrect calibration can distort profiles and introduce noise in time of arrival (TOA) data. In a previous article we presented a new method for calibrating pulsar observations conducted with the Nançay decimetric Radio Telescope (NRT), which significantly improved NRT polarimetric measurements and pulsar timing quality for data taken after this method was developed, in November 2019.
Aims. Results from the above-mentioned study hinted at a dependence of the polarimetric response of the NRT on the observed direction. We therefore investigated this potential dependence, since unaccounted variations in the instrumental response could degrade polarimetric measurements. Additionally, our aim was to develop a method for properly calibrating NRT pulsar observations conducted before November 2019.
Methods. We conducted three series of observations of bright pulsars over wide declination ranges, in a special observation mode in which the feed horn rotates by 180∘ degrees across the observation; this enabled us to determine the full polarimetric response of the NRT while modeling potential variations in calibration parameters with hour angle and declination. In addition, we developed a method that uses the measurement equation template matching (METM) technique to improve the calibration of pre-November 2019 data.
Results. From the analysis of the series of observations of bright pulsars with horn rotation, we found that the polarimetric response of the NRT does not appear to vary with hour angle or declination. On the other hand, the new METM-based calibration method appears to significantly improve the calibration of pre-November 2019 data. By analyzing NRT data on a selection of millisecond pulsars, we found that the new polarimetric profiles are more homogeneous, they generally have higher signal-to-noise ratios, and found that the TOA data for these MSPs are more accurate and contain lower levels of noise, especially when combining the new calibration method with the matrix template matching (MTM) method for extracting TOAs from pulsar observations.
@@
Pulsars are rotation-powered neutron stars that emit beams of electromagnetic radiation. These beams are swept across the sky with the rotation of the stars, so that distant observers whose lines of sight are crossed by the beams might detect periodic pulses. The pulsar timing technique consists in modeling the rotation of pulsars by measuring the times of arrival (TOAs) of the pulses detected at a telescope and comparing them to the predictions from a timing model. Differences between measured and predicted TOAs, called timing residuals, contain information about the pulsars’ spin properties, orbital properties (for pulsars in multiple systems), astrometric properties, or information about the interstellar medium or their immediate environment (see, e.g., Chapter 2 of Lorimer & Kramer, 2004, for a complete description). Applications of pulsar timing are numerous and diverse: for example, tests of general relativity (GR) in the strong-field regime (Kramer et al., 2021); searches for exoplanets orbiting pulsars (Niţu et al., 2024; Voisin et al., 2025); measurements of pulsar masses and constraints on the equations of state of neutron star masses using these measurements (e.g., Özel & Freire, 2016); or searches for low-frequency gravitational waves using pulsar timing arrays (PTAs; see FERMI-LAT Collaboration et al., 2022; Agazie et al., 2023; Reardon et al., 2023; EPTA Collaboration et al., 2023b, for recent results from the different PTA collaborations).
@@
In standard radio pulsar timing analyses, TOAs for a given pulsar are determined by comparing radio observations to a standard profile, assumed to be a noise-free representation of the pulsar’s intrinsic pulse profile shape. The standard profiles can be used to determine TOAs for arbitrarily large fractions of the recorded frequency bandwidths, assuming that the pulse profile does not evolve with frequency within these frequency intervals. Alternatively, the standard profiles can encode frequency evolution of their component(s), leading to more accurate TOAs in the case of large bandwidth observations (the latter wide-band template matching technique is for instance implemented in the PulsePortraiture software library, see Pennucci et al., 2014). Independently of the method used for constructing the template profiles, standard pulsar timing analyses crucially rely on the assumption that the template profiles are faithful representations of the pulsar pulse profiles. Reciprocally, individual observations of a given pulsar must resemble the standard profile and not vary, otherwise the cross-correlation between the individual observations and the standard profile will lead to inaccurate TOAs and introduce systematic timing errors in the TOAs.
@@
If not properly accounted for, pulse profile variations can cause individual observations of a pulsar to differ from the chosen standard profile. In some cases, these variations originate from the pulsar itself, e.g., from pulse jitter, moding, intermittency, or relativistic spin precession (see Padmanabh et al., 2021, and references therein). Profile variations can also result from instrumental effects. In particular, inaccurate polarization calibration can distort recorded pulse profiles and in turn introduce noise in TOA datasets (see for instance van Straten, 2013; Foster et al., 2015; Guillemot et al., 2023; Rogers et al., 2024). Such timing noise resulting from instrumental artifacts can degrade pulsar timing analyses, bias parameter measurements, or mimic signals expected from stochastic gravitational wave background emission (Lentati et al., 2016).
@@
Accurate polarization calibration is therefore crucial for any pulsar timing studies, and for studying pulsar polarization properties. Recently, Dey et al. (2024) investigated different methods for calibrating pulsar observations done with the Green Bank Telescope, and found that the simplest method, the ideal feed assumption (hereafter IFA; Manchester et al., 2013) which assumes perfectly linear and orthogonal polarization feeds, led to better pulsar timing results than the more sophisticated measurement equation modeling (MEM, van Straten, 2004) and measurement equation template matching (METM, van Straten, 2013) methods in which polarimetric responses are respectively determined from the analysis of strongly linearly polarized pulsars observed over wide ranges of parallactic angles, or by comparing observations of a given reference pulsar with a well-calibrated template profile for the same pulsar. Dey et al. (2024) attributed the poorer performance of the other tested methods to potential instabilities of the local polarization reference source, a noise diode coupled to the receiver, and time instabilities in the pulse profile of the pulsar used as reference for the METM method. The IFA, MEM, and METM calibration methods are described in more detail later in this paper. Rogers et al. (2024), on the other hand, compared polarization calibration methods using pulsar data from the Murriyang 64m CSIRO Parkes radio telescope, and found that the METM method significantly improved pulsar timing quality over the IFA calibration scheme. Furthermore, Rogers et al. (2024) found that combining the METM data calibration method with the matrix template matching (MTM) technique (van Straten, 2006), which uses all four Stokes components of the pulse profile and template profile when forming TOAs to compensate for residual inconsistencies in the polarimetric profiles, leads to significantly improved pulsar timing residuals. In their conclusions, Rogers et al. (2024) recommended that MTM be used for determining TOAs in PTA experiments. Finally, in previous work (see Guillemot et al., 2023) we significantly improved the polarization calibration of pulsar observations made with the Nançay decimetric Radio Telescope (NRT) in France by determining full polarimetric responses of the instrument from the analysis of observations of a highly linearly polarized pulsar in a special observing mode where the feed horn rotates across the observation, and by employing the MEM calibration model (details of this analysis are given in Sect. 2). The conclusion from this work is that the MEM-based calibration scheme outperforms IFA calibration for the NRT, with much improved polarimetric measurements, contrasting with the conclusions of Dey et al. (2024). In addition, similarly to Rogers et al. (2024) we also found that determining TOAs using the MTM technique significantly improved the NRT pulsar timing quality, which also suggested that this TOA extraction technique should be used in high-precision pulsar timing studies.
@@
In this article, we present new work done following Guillemot et al. (2023) to further improve the polarization calibration of pulsar data taken with the NRT. The results from the observations of a bright highly linearly polarized pulsar with feed horn rotation presented in Guillemot et al. (2023) hinted at an hour angle-dependence of the response of the NRT. We therefore conducted an analysis aimed at determining whether the instrumental response of the NRT depends on the observed direction, and more specifically on the hour angle and declination of the observed pulsar. In Sect. 3 we examine the possibility of properly calibrating NRT data taken before the first observations, in November 2019, of bright highly polarized pulsars with horn rotation. This work, that used the METM calibration technique, led to the construction of a time- and frequency-dependent calibration model that we applied to observations of millisecond pulsars (MSPs) carried out with the NRT. In Sect. 4 we present the new polarimetric profiles for these MSPs as well as an assessment on pulsar timing quality with the calibration scheme. We finally present a summary of our work in Sect. 5.
2 Polarimetric response of the NRT: Possible dependence on the observed direction
@@
The NRT is a meridian transit-type telescope of the Kraus/Ohio State design (see, for instance, Kraus, 1960, 1966) equivalent to a parabolic dish with a diameter of approximately 94m, and capable of observing objects with declinations above . Since August 2011, NRT pulsar observations are conducted with the NUPPI backend which uses Graphics Processing Units (GPUs) to coherently dedisperse and fold the dual linear polarization signals from the receiver in real time, over a total bandwidth of 512 MHz (see, e.g., Desvignes et al., 2011; Cognard et al., 2013, for additional details on NUPPI). Pulsar observations with the NRT are primarily conducted at the central frequency of 1484 MHz. Further details on the NRT and on pulsar observations with the NUPPI backend can be found in Guillemot et al. (2023). In this new article, we consider pulsar observations made with the NUPPI data at 1.4 GHz, leaving analyses of observations at other frequencies or observations conducted with the previous backend, the Berkeley-Orléans-Nançay (BON) pulsar backend, for future work. For all NUPPI data analyses presented in this new article we made use of the PSRCHIVE software library (Hotan et al., 2004), and cleaned the data of radio-frequency interference (RFI) using the Surgical method of the CoastGuard pulsar data analysis library (Lazarus et al., 2016).
@@
In the first article of this series (Guillemot et al., 2023), we presented an improved procedure for calibrating pulsar observations conducted with the NRT, developed in late 2019. Owing to its design, the NRT can track objects for, typically, one to two hours around transit (the maximum allowed tracking time depends on the observed declination). As a consequence, the NRT cannot sample wide ranges of parallactic angles when observing a given object. This limitation, in principle, prevents measurements of the full polarimetric response of the telescope with long observations of bright, highly polarized pulsars over wide ranges of parallactic angles (see, e.g., Stinebring et al., 1984; Johnston, 2002; van Straten, 2004, for examples of such experiments). Nevertheless, in November 2019 we circumvented this limitation by beginning to carry out regular observations of the bright, highly linearly polarized pulsar J07422822 (also known as B074028, see Bonsignori-Facondi et al., 1973), in a non-standard observation mode in which the feed horn rotates by over the course of the 1 h observation. During these observations, the feed horn rotation effectively mimics wide parallactic angle variations, enabling much wider ranges of pulsar polarization angles to be sampled over the hour of observation. In turn, this enabled us to determine the full polarimetric response of the telescope. The calibration solutions obtained from these rotating horn observations enabled us to obtain vastly improved polarimetric results for the selection of millisecond pulsars (MSPs) that was used in our tests, with increased signal-to-noise ratios (S/N) in most cases and polarimetric profiles that were consistent with published results, unlike those obtained with the previously used, simplistic calibration method.
@@
We began by investigating our analysis codes, and found an issue with the calculation of the time-varying parallactic for the NRT, in the analysis that led to the results presented in Fig. 9 of Guillemot et al. (2023). In the case of the NRT, the parallactic angle is given by
@@
where denotes the declination of the observed source, HA is the hour angle of the source at a given time, and is the parallactic angle. The above expression is the same as the one given in Guillemot et al. (2023), and was the expression implemented in PSRCHIVE. However, we found that, due to the conventions used in PSRCHIVE for the signs and orientations of angles, the parallactic angle calculated using Eq. 1 needed to be used with the opposite sign in the software. We found that correcting for the sign of for NRT observations in PSRCHIVE significantly improved the results of the analysis of rotating horn observations. In Fig. 1 we show measured and modeled Stokes parameters for an observation of J07422822 with horn rotation conducted on MJD 59368 (3 June 2021), before and after correction. As can be seen from the right panel of Fig. 1, measured and modeled values of the Stokes parameter in the first half of the observation are much more consistent. The fit results shown in the left panel of Fig. 1 (i.e., before correction) give a reduced statistic of 2.25, while those shown in the right panel give 1.73, indicating a better agreement between the data and the model. The analysis of all other rotating horn observations of J07422822 also found a better agreement between the data and the model, after the sign correction. In addition to this correction we made other improvements to the analysis pipeline. First, variations in the ionospheric Rotation Measure (RM) are now taken into account, using RMextract (Mevius, 2018) to determine ionospheric RM values for each sub-integration. Second, we now make use the “-Q” option of pcm to indicate that the reference noise diode is coupled after the frontend, as is indeed the case for our observations. We note, however, that the latter two modifications of the analysis pipeline compared to Guillemot et al. (2023) resulted in minor improvements to the analysis results.
@@
Beyond the correction made to the calculation of NRT parallactic angles in PSRCHIVE that significantly improved the agreement between measured and modeled Stokes parameters for rotating horn observations, the initial goal of the study was to test for potential dependencies of the response of the NRT on observation parameters. As was noted in Guillemot et al. (2023), the absolute gain of the NRT is known to vary with the declination of the observed object (see, e.g., Theureau et al., 2005). To test for a dependence of calibration parameters on declination, and also test for a dependence on hour angle (e.g., resulting from the unusual design of the telescope), we therefore conducted observations of pulsars other than J07422822 in the special observation mode. We conducted a first series of rotating horn observations of pulsars at different declinations between MJDs 59368 and 59373 (3 and 8 June 2021). During this first series of observations, we sampled pulsar declinations between (J07422822) and (J0454+5543, a.k.a. B0450+55, Damashek et al., 1978). The other pulsars observed during this session were PSRs J0826+2637 (B0823+26, Craft et al., 1968), J0953+0755 (B0950+08, Pilkington et al., 1968), J1239+2453 (B1237+25, Lang, 1969), J17051906 (B170219, Manchester et al., 1978), and J2113+4644 (B2111+46, Davies & Large, 1970). The observations were conducted close in time (in this case, within five days), so that we could assume the instrumental response to be similar at all epochs. Table 1 lists the start epochs, declinations, durations, and start and end hour angles of these observations. Figs. 2 and 3 show the results of the MEM analysis of the individual observations and preceding observations of the reference noise diode. As in Fig. 1, left (resp., right) panels show the results obtained before (resp., after) correction of the sign of the NRT parallactic angle. Interestingly, inconsistencies between pulsars for some parameters (particularly, the differential gain , the feed orientation parameter and the Stokes parameter of the reference signal, i.e., parameters that constrain the polarization position angle, which strongly depends on the parallactic angle correction), appear to be significantly reduced by the correction. After correction, calibration and reference signal parameters seem consistent between pulsar observations, with the exception of the absolute gain , known to vary with declination.
@@
Strong RFI conditions can corrupt some of the observations, e.g., leading to entire frequency channels being blanked by RFI excision in these observations. If other observations are less affected then the analysis will determine the polarimetric response in more frequency channels than that of the individual, RFI-corrupted observations.
@@
Prior to this work, the pcm tool of PSRCHIVE did not allow multiple pulsar observations to be analyzed jointly. Therefore, in order to be able to conduct analyses of this type, we added new functionalities to pcm222Currently available under the “npsr-pcm” branch of PSRCHIVE, see git://git.code.sf.net/p/psrchive/code..
@@
In the above expressions, represents the 2D polynomial, is the degree of the polynomial in and is the degree of the polynomial in , and the terms represent coefficients to be fitted for. In our analysis, and respectively corresponded to the pulsar declination , and to the hour angle HA at a given time. We tested polynomial orders ranging from 0 (i.e., constant values) to 3 in HA, and ranging from 0 to 6 in declination. We used the same polynomial orders for both the differential gain and phase functions, to limit the number of polynomial order combinations to be tested. In the rest of the article we denote the different analyses we performed using the following nomenclature: “A/B”, meaning that the 2D polynomials representing the differential gain and phase were of order A in declination, and of order B in hour angle.